How can you find the number of seats in the succeding rows?


1
Expert's answer
2020-10-08T17:00:09-0400

By using Arithmetic progression.

For example - first row has "15" seats, second row has "18" seats and third row has "21" seats and so on.

So we can see a constant rise in number of seats called common difference that is "d= 3."

Seats in first row "a_1= 15"

We can find seats in "n^{th}" row by using the formula "a_n= a_1 +(n-1) d"

To find seats in "5th" row, "n= 5."

So seats in fifth row are

"a_5= a_1+ (n-1)d= 15+(5-1)3=27"